Update from Stony Brook Stony Brook Gang
Two More Shots • Seem to show Good QE in local tests: • https://www.phenix.bnl.gov/WWW/p/draft/jkamin/browse1_qe.html • Chicklet mystery: • Expect 150 pA current. • Saw 0.7 pA??? • Mishandling cause factor of 200 QE drop?? • Still have another Chicklet, same shot.
Pump Woes • Main Scroll pump suffered failure. • Rebuilt one on order. • Local fix to existing one as well. • Used “little guy” as main pump. • Revealed additional troubles: • Cryo gate valve leaks! • Too large gas load during cool-down. • Burden on turbo during pumpdown. • Rich L and Bill will service it this week.
Bad hole on Gain Test Chamber. • One LEMO hole is bigger than others! • LEMO does not engage. • Will weld plug over the hole to seal.
No shot this week. • Concentrate on curing the inconsistent performance of the cryo pump: • Cryo Gate Valve CERTAINLY leaks. • ALL observations could be consistent with a single failure: Cryo Main Gate Valve. • Time for minor mods to Evap chamber. • Next “weekly shot” next week.